OpenStudy (anonymous):

What is the factored form of 2x^2-8x-42 ?

OpenStudy (anonymous):

2(x-7)(x+3)

OpenStudy (anonymous):

(2)(x+3)(x-7)

OpenStudy (anonymous):

2x^2-8x-42 =2x^2-14x+6x-42 (middel factor theorem) =2x(x-7)+6(x-7) =(x-7)(2x+6)